. A company has 6 senior and 5 junior officers. An ad hoc legislative

   committee is to be formed. In how many ways can a 5-officer committee be

   formed so that it is composed of:

   a) Any five officers (5 pts.)

   b) 3 senior officers and 2 junior officers (5 pts.)

(PLEASE TYPE THE ANSEWERS IN THE COMMENT BOX. PLEASE DON'T WRITE THE ANSWER ON A SHEET OF PAPER) THANKS AND I WILL LEAVE GREAT FEEDBACK
Answer #1

Toal officers = 6 + 5 = 11

a) Number of ways to select a committee of any 5 officers = 11C5 = 11! / (5! * (11 - 5)!) = 462

b) Number of ways to select a committee of 3 senior officers and 2 junior officers = 6C3 * 5C2 = 20 * 10 = 200